OpenStudy (anonymous):

Calculus Q: Assume f and g are differentiable functions with h(x) = f(g(x)). Suppose the equation of the line tangent to the graph of g at the point (4,7) is y = 3x-5 and the equation of the line tangent to the graph of f at (7,9) is y = -2x + 23. (a) Calculate h(4) and h'(4). (b) Determine an equation of the line tangent to the graph of h at the point on the graph where x = 4.

OpenStudy (anonymous):

The slope of the tangent line to \(g\) at \(x=4\) is \(3\), i.e. \(g'(4)=3\). Similarly, you see that \(f'(7)=-2\). So, if \(h(x)=f(g(x))\), then \(h(4)=f(g(4))=f(7)=9\). Applying the chain rule, you have \(h'(x)=f'(g(x))\cdot g'(x)\), so \(h'(4)=f'(g(4))\cdot g'(4)=f'(7)g'(4)=-2(3)=-6\). You can use this info to find the equation of the tangent line to \(h\) when \(x=4\).

OpenStudy (anonymous):

so basicall this is what I did: g(4) = 7 g'(4) = 12 f(7) = 9 f'(7) = -14 h(x)=f(g(x)) h'(x)=f'(g(x))*g'(x) h(4) = f(g(4)) = f(7) = 9 h'(4) = f(g(4))*g'(4) = f(7) * 12 = 84

OpenStudy (anonymous):

but the slope is way too big!!

OpenStudy (anonymous):

How did you get that value for g'(4)? and f'(7)?

OpenStudy (anonymous):

ahhh. ok so the function is y=3x-5 and to get g'4 I took the 3 and subbed in 4 for x to get 12

OpenStudy (anonymous):

and to get g(4) ...i just subbed in 4 for x and got 7.

OpenStudy (anonymous):

my derivative for g'4 is wrong...how did you get your value?

OpenStudy (anonymous):

The derivative of g at x=4 is the same as the slope of the tangent line at x=4. You need only check the slope of the tangent line.

OpenStudy (anonymous):

eeek thank-you! and would the equation be: y-9=-6(x-4)

OpenStudy (anonymous):

You're welcome, and yes.